# Still More Shades of Null: A Benchmark for Responsible Missing Value Imputation

### v1

@@ Abstract.
We present Shades-of-Null, a benchmark for responsible missing value imputation. Our benchmark includes state-of-the-art imputation techniques, and embeds them into the machine learning development lifecycle. We model realistic missingness scenarios that go beyond Rubin’s classic Missing Completely at Random (MCAR), Missing At Random (MAR) and Missing Not At Random (MNAR), to include multi-mechanism missingness (when different missingness patterns co-exist in the data) and missingness shift (when the missingness mechanism changes between training and test). Another key novelty of our work is that we evaluate imputers holistically, based on the predictive performance, fairness and stability of the models that are trained and tested on the data they produce.
We use Shades-of-Null to conduct a large-scale empirical study involving 20,952 experimental pipelines, and find that, while there is no single best-performing imputation approach for all missingness types, interesting performance patterns do emerge when comparing imputer performance in simpler vs. more complex missingness scenarios. Further, while predictive performance, fairness and stability can be seen as orthogonal, we identify trade-offs among them that arise due to the combination of missingness scenario, the choice of an imputer, and the architecture of the model trained on the data post-imputation. We make Shades-of-Null publicly available, and hope to enable researchers to comprehensively and rigorously evaluate new missing value imputation methods on a wide range of evaluation metrics, in plausible and socially meaningful missingness scenarios.
@@ 1. Introduction
Specifically in data management, debates have been raging since the very inception of the field regarding whether — and how — to handle missing values, see  Date (1984) for a classic example. One difficulty, at the operational level, is that, while missing values are typically denoted with the token null, that may not always be the case. There could be hidden missing values in the data, like when ‘AL’ is selected as the state on a job application because it’s the first item in the pull-down list. Another difficulty, at the semantic level, is that a null can take on many meanings: it may signify that the value is unknown, inapplicable, intentionally withheld, etc.
In this paper, we do not concern ourselves with this semantic debate, and we also do not consider hidden missing values. Rather, we focus our attention on a very specific case. We are given a dataset $X$ (a single relation), in which the values of some features are missing, denoted by the marker null. The meaning of null is that the feature does take on a value in the real world (i.e., it is applicable) but that value is unobserved in $X$. Further, we intend to use $X$ in a machine learning (ML) setting: to train a model or to make predictions about the tuples in $X$ using a model that has already been trained and deployed. This model cannot deal with null values directly, and so these values must be imputed (filled in) as part of data pre-processing.
@@
Rubin’s framework has formed the basis of a large body of technical work on missing value imputation (MVI), which has been covered in several excellent and comprehensive surveys (Abdelaal et al., 2023; Hasan et al., 2021; Le Quy et al., 2022; Mena et al., 2022; Pintz et al., 2022; Emmanuel et al., 2021; Peherstorfer et al., 2018; Lin and Tsai, 2020; Pigott, 2010; Zhou et al., 2024; Alabadla et al., 2022; García Laencina et al., 2010; Joel and Doorsamy, 2022; Ilyas and Chu, 2019). Broadly speaking, MVI approaches fall into several categories: (1) statistical, such as imputing missing observations with the median or mode (Schafer and Graham, 2002); (2) learning-based impute-then-classify, which use the observed data to iteratively impute and learn missing values, using k-nearest neighbors (Batista and Monard, 2003), clustering (Gajawada and Toshniwal, 2012), decision trees (Twala, 2009), or ensembles (Stekhoven and Bühlmann, 2012); and (3) joint data cleaning and model training (Karlaš et al., 2020; Krishnan et al., 2017, 2016), based on Rubin’s seminal multiple imputation framework (Rubin, 1987).
@@
Despite an abundance of MVI techniques using different model architectures being proposed each year, making meaningful progress on learning from incomplete data ultimately comes down to how realistically we model missingness: our proposed methods are only as good as the evaluation scenarios we rigorously test them on. If our evaluation scenarios do not realistically capture the kinds of missingness that we encounter in practice, then our data cleaning methods are unlikely to be effective. Further, newly proposed MVI techniques are often only evaluated on the correctness of the imputation (imputation quality) and the correctness of the predictive model (accuracy or F1). However, in recent years, the scientific community is starting to uncover trade-offs between MVI efficacy and other normatively important dimensions of model performance, including algorithmic fairness (Guha et al., 2024; Feng, 2023), and stability and robustness (Zhou et al., 2024), beyond accuracy. We elaborate on this next.
@@
Consider again the job applicant screening example, and suppose that the applicants’ gender and age are among the features. Female job applicants who suspect that they have been receiving lower salaries than their male colleagues, despite having comparable qualifications and job experience, may withhold salary information more frequently then men, hoping to narrow the gender wage gap in their next position. As a result there will be more missing salary values for women, and the probability that the value is missing depends on the observed covariate, gender. This situation is consistent with MAR missingness, and it points to pre-existing bias — a situation where data reflects a history of societal discrimination (Friedman and Nissenbaum, 1996). For our final example of pre-existing bias, suppose that disability status is among the features. Job applicants who live with a disability are more likely to omit a value for this feature. Assuming that this feature does not correlate with the other features, this situation is consistent with MNAR, with missingness itself serving as a proxy for disadvantage.
@@
Our work is (1) novel: to the best of our knowledge, the settings of multi-mechanism missingness and missingness shifts have not been empirically studied before; (2) comprehensive: we evaluate a suite of 8 MVI techniques on 6 benchmark datasets using 5 model types, running a total of 20,952 pipelines, and is the first study of such scale in the missing data domain to the best of our knowledge; (3) normatively grounded: we focus on decision-making contexts such as lending, admissions, and healthcare, where missingness is socially salient. Mitigating social harm such as algorithmic discrimination is a leading concern in these domains, and we evaluate the impact of MVI approaches on downstream model fairness and stability (which have been understudied in the context of missing data), in addition to classically studied imputation quality and model correctness metrics.
As an additional engineering contribution, while developing the Shades-of-Null benchmark, we improved several state-of-the-art MVI techniques, see Appendix A.2 for details.
We make Shades-of-Null publicly available [^1] https://github.com/FalaahArifKhan/data-cleaning-stability and hope to enable researchers to comprehensively and rigorously evaluate new missing value imputation methods on a wide range of evaluation metrics, in plausible and socially meaningful missingness scenarios.
@@ 2. Related Work
Missing value imputation (MVI) techniques
Learning-based approaches have become increasingly popular, and include k-Nearest Neighbors, Decision Trees, Support Vector Machines, clustering, and ensembles (Batista and Monard, 2003; Lin and Tsai, 2020; Hasan et al., 2021).  Zhou et al. (2024) and Liu et al. (2023) review Deep Learning-based approaches (Variational Auto-Encoders and GANs) and representation learning (Graph Neural Networks and diffusion-based methods). Multiple imputation (Schafer and Graham, 2002; Zhou et al., 2024) and Expectation Maximization (Pigott, 2010) are also influential, but too computationally expensive to be popular in practice (Alabadla et al., 2022). According to (Feng, 2023; Le Morvan et al., 2021), impute-then-classify is suboptimal compared to multiple imputation because missingness itself can be informative and is lost before model training.
@@
Empirical studies have been overwhelmingly focused on medical datasets, and only evaluate missingness under MCAR (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008). Further, most proposed methods only perform direct evaluation, using metrics such as MAE, MSE, RMSE, and AUC (Hasan et al., 2021; Joel and Doorsamy, 2022). Very few studies use both direct and indirect evaluation, and even studies that include indirect evaluation only evaluate predictive accuracy (Lin and Tsai, 2020). This leaves a significant research gap, specially for desiderata such as stability and self-consistency, which have recently been shown to be highly sensitive to data-centric interventions (Black and Fredrikson, 2021; Black et al., 2022; Cooper et al., 2023), and unfairness, which can be worsened by improperly chosen data cleaning steps (Guha et al., 2024). Additionally, the performance of MVI techniques under multi-mechanism missingness (Zhou et al., 2024) and missingness shifts (Zhou et al., 2023) remains unexplored, despite these conditions being more likely to occur in practice due to distribution shifts in production deployments (Graham, 2009).
Notably, and worryingly, overwhelming evidence in the literature indicates that there is no single “best-performing” MVI approach on accuracy  (Hasan et al., 2021; Lin and Tsai, 2020; Shadbahr et al., 2023; García Laencina et al., 2010; Farhangfar et al., 2008), and that model performance (narrowly measured based on ‘correctness’ thus far) depends on dataset characteristics (including size and correlation between variables) (Alabadla et al., 2022) and missingness rates in the train and test sets (Lin and Tsai, 2020; Shadbahr et al., 2023).
@@
There has been some recent interest in studying the social harm that can come from poorly chosen missing value imputation techniques (Caton et al., 2022; Zhang and Long, 2022, 2021a, 2021b; Jeong et al., 2022; Wang and Singh, 2021; Martínez-Plumed et al., 2019; Feng et al., 2024). Most empirical studies (Wang and Singh, 2021; Zhang and Long, 2021b, 2022; Caton et al., 2022) have worked with the COMPAS and Adult datasets, the latter of which has been “retired” from community use due to issues with provenance (Ding et al., 2021). These datasets do not contain naturally-occurring missing values, and experimental studies employ randomly-generated missingness: usually by sampling or using a fixed set of columns, and randomly picking rows in which to replace values with null. We critique this approach, since detecting and mitigating unfairness requires broader socio-technical thinking, such as having higher rates of missingness for minority groups and in features that are highly correlated with sensitive attributes (called proxy variables in the fairness literature) (Caton et al., 2022).
@@
Zhang and Long (2021b) evaluate MVI methods on imputation fairness, measured as the imputation accuracy difference. They find that (1) higher missingness disparity leads to higher imputation unfairness; (2) higher rates of missingness lead to higher imputation unfairness; (3) more imbalance in the data (or larger disparity in the sample sizes across groups) leads to higher imputation unfairness, since there is an imbalance of observed information; (4) the same imputation method has different prediction fairness under different missingness mechanisms. Their evaluation is limited to randomly generated null values in COMPAS. We extend their analysis of fairness trade-offs to additional datasets and missingness scenarios.
Zhang and Long (2022) introduce the notion of imputation fairness risk as the disparity in imputation quality for different demographic groups, and provide bounds on the fairness risk of “correctly specified” imputation methods. While there is very little theoretical work in this space, and this is a commendable contribution, we call into question the broader implications of this result: it is well-known that imputation quality metrics do not fully capture downstream model performance (Zhang and Long, 2021a). Put differently, a classifier built on poor imputation quality can still give good downstream performance (Shadbahr et al., 2023). With this in mind, we ask the question: Does minimizing imputation unfairness minimize model unfairness? Our empirical findings indicate that they do not, as is discussed in Sections 5.4.
Missingness and stability
We are not aware of any work investigating the effect of missing value imputation on model stability.
@@ 3.1. Methodology for Simulating Missingness
In this work, we start with datasets in which there are no missing values, and then simulate missingness. We make this choice to be able to evaluate the effectiveness of missing value imputation scenarios with respect to the ground truth. Our methodology for simulating missingness is based on evaluation scenarios, defined by the missingness mechanism during train and test, shown in Table 1: (1) single-mechanism missingness, injected similarly into train and test (S1 - S3); (2) single-mechanism missingness, injected differently into train and test (missingness shift) (S4 - S9); and (3) missingness is mixed, to include all three missingness mechanisms, and is injected similarly into train and test (S10).
@@
For each dataset in our benchmark (see Section 3.5), we devised dataset-specific missingness scenarios corresponding to the three missingness mechanisms (MCAR, MAR, MNAR). These scenarios were crafted to be realistic, and were guided by the correlation coefficients between all covariates and the target variable, as well as by feature importance derived using scikit-learn’s built-in functionality, following (Joel and Doorsamy, 2022; Martínez-Plumed et al., 2019). Furthermore, we selected columns that may plausibly contain missingness. For example, blood pressure or cholestrol in diabetes is expected to always be observed, but family history or physical activity might be omitted or withheld. Table 3 presents missingness scenarios for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2. Details for other datasets are deferred to the Appendix B.2.
@@ 3.2. Missing Value Imputation (MVI) Techniques
As discussed in Section 2, there is a large number of competitive MVI techniques in the literature. For our study, we picked 8 techniques, from 5 broad categories in the taxonomy of Emmanuel et al. (Emmanuel et al., 2021), namely: (1) baseline: deletion; (2) statistical techniques: median-mode and median-dummy; (3) machine learning-based techniques: miss-forest (Stekhoven and Bühlmann, 2011) and clustering (Gajawada and Toshniwal, 2012); (4) deep learning-based techniques: datawig (Biessmann et al., 2019) and auto-ml (Jäger et al., 2021); and (5) multiple imputation: boostclean (Krishnan et al., 2017). See Appendix A.1 for details. We limit our analysis to one joint-cleaning-and-training approach, boostclean (Krishnan et al., 2017), and leave the exploration of other methods, such as ActiveClean (Krishnan et al., 2016) and CPClean (Karlaš et al., 2020) to future work.

### v2

# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

@@ Abstract.
Data missingness is a practical challenge of sustained interest to the scientific community. In this paper, we present Shades-of-Null, an evaluation suite for responsible missing value imputation. Our work is novel in two ways (i) we model realistic and socially-salient missingness scenarios that go beyond Rubin’s classic Missing Completely at Random (MCAR), Missing At Random (MAR) and Missing Not At Random (MNAR) settings, to include multi-mechanism missingness (when different missingness patterns co-exist in the data) and missingness shift (when the missingness mechanism changes between training and test) (ii) we evaluate imputers holistically, based on imputation quality, as well as on the predictive performance, fairness and stability of the models that are trained and tested on the data post-imputation.
We use Shades-of-Null to conduct a large-scale empirical study involving 23,940 experimental pipelines, and find that while there is no single best-performing imputation approach for all missingness types, interesting trade-offs arise between predictive performance, fairness and stability, based on the combination of missingness scenario, imputer choice, and the architecture of the predictive model. We make Shades-of-Null publicly available, to enable researchers to rigorously evaluate missing value imputation methods on a wide range of metrics in plausible and socially meaningful scenarios.
@@ 1. Introduction
Specifically in data management, debates have been raging since the very inception of the field regarding whether — and how — to handle missing values, see  Date (1984) for a classic example. One difficulty, at the operational level, is that, while missing values are typically denoted with the token null, that may not always be the case. There could be hidden missing values in the data, like when ‘AL’ is selected as the state on a job application because it’s the first item in the drop-down list. Another difficulty, at the semantic level, is that a null can take on many meanings: it may signify that the value is unknown, inapplicable, intentionally withheld, etc.
In this paper, we do not concern ourselves with this semantic debate, and we also do not consider hidden missing values (Qahtan et al., 2018). Rather, we focus our attention on a very specific case. We are given a dataset $X$ (a single relation), in which the values of some features are missing, denoted by the marker null. The meaning of null is that the feature does take on a value in the real world (i.e., it is applicable) but that value is unobserved in $X$. Further, we intend to use $X$ in a machine learning (ML) setting: to train a model or to make predictions about the tuples in $X$ using a model that has already been trained and deployed. This model cannot deal with null values directly, and so these values must be imputed (filled in) as part of data pre-processing.
@@
Rubin’s framework has formed the basis of a large body of technical work on missing value imputation, which has been covered in several excellent and comprehensive surveys (Abdelaal et al., 2023; Hasan et al., 2021; Le Quy et al., 2022; Mena et al., 2022; Pintz et al., 2022; Emmanuel et al., 2021; Peherstorfer et al., 2018; Lin and Tsai, 2020; Pigott, 2010; Zhou et al., 2024; Alabadla et al., 2022; García Laencina et al., 2010; Joel and Doorsamy, 2022; Ilyas and Chu, 2019). Broadly speaking, MVI approaches fall into several categories: (1) statistical, such as imputing missing observations with the median or mode (Schafer and Graham, 2002); (2) learning-based impute-then-classify, which use the observed data to iteratively learn and impute missing values, using k-nearest neighbors (Batista and Monard, 2003), clustering (Gajawada and Toshniwal, 2012), decision trees (Twala, 2009), or ensembles (Stekhoven and Bühlmann, 2012); and (3) joint data cleaning and model training (Karlaš et al., 2020; Krishnan et al., 2017, 2016), based on Rubin’s seminal multiple imputation framework (Rubin, 1987).
@@
Despite an abundance of MVI techniques using different model architectures being proposed each year, making meaningful progress on learning from incomplete data ultimately comes down to how realistically we model missingness: our proposed methods are only as good as the evaluation scenarios we rigorously test them on. If our evaluation scenarios do not realistically capture the kinds of missingness that we encounter in practice, then our data cleaning methods are unlikely to be effective. Further, newly proposed MVI techniques are often only evaluated on the correctness of the imputation (imputation quality) and the correctness of the predictive model (accuracy or F1). However, in recent years, the scientific community is starting to uncover trade-offs between MVI efficacy and other normatively important dimensions of model performance, including algorithmic fairness (Guha et al., 2024; Feng, 2023), and stability and robustness (Zhou et al., 2024), beyond accuracy.
@@
Consider again the job applicant screening example, and suppose that the applicants’ gender and age are among the features. Female job applicants who suspect that they have been receiving lower salaries than their male colleagues, despite having comparable qualifications and job experience, may withhold salary information more frequently then men, hoping to narrow the gender wage gap in their next position. As a result there will be more missing salary values for women, and the probability that the value is missing depends on the observed covariate, gender. This situation is consistent with MAR missingness, and it points to pre-existing bias — a situation where data reflects a history of societal discrimination (Friedman and Nissenbaum, 1996). For our final example, suppose that disability status is among the features. Job applicants who live with a disability are more likely to omit a value for this feature. Assuming that this feature does not correlate with the other features, this situation is consistent with MNAR, with missingness itself serving as a proxy for disadvantage.
@@
Our work is (1) novel: to the best of our knowledge, the settings of multi-mechanism missingness and missingness shifts have not been empirically studied before; (2) comprehensive: we evaluate a suite of 8 MVI techniques on 7 benchmark datasets using 6 model types, running a total of 23,940 pipelines, and is the first study of such scale in the missing data domain, to the best of our knowledge; (3) normatively grounded: we focus on decision-making contexts such as lending, hiring, and healthcare, where missingness is socially salient. Mitigating social harm such as algorithmic discrimination is a leading concern in these domains (Barocas and Selbst, 2016), and we evaluate the impact of MVI approaches on downstream model fairness and stability (which have been understudied in the context of missing data), in addition to classically studied imputation quality and model correctness metrics.
While developing the Shades-of-Null evaluation suite, we found and fixed several bugs in existing MVI implementations, including data leakage and omitted hyperparameter tuning. See Appendix A.2 for details.
We make Shades-of-Null publicly available [^1] https://github.com/FalaahArifKhan/data-cleaning-stability and hope to enable researchers to comprehensively and rigorously evaluate new MVI methods on a wide range of evaluation metrics, under plausible and socially meaningful missingness scenarios.
@@ 2. Related Work
Missing value imputation techniques
Learning-based approaches have become increasingly popular, and include k-nearest neighbors, decision trees, support vector machines, clustering, and ensembles (Batista and Monard, 2003; Lin and Tsai, 2020; Hasan et al., 2021).  Zhou et al. (2024) and Liu et al. (2023) review deep learning-based approaches (variational auto-encoders and generative adversarial networks) and representation learning (graph neural networks and diffusion-based methods). Multiple imputation (Schafer and Graham, 2002; Zhou et al., 2024) and expectation maximization (Pigott, 2010; Wang et al., 2024) are also influential, but too computationally expensive to be popular in practice (Alabadla et al., 2022).
@@
Empirical studies have been overwhelmingly focused on medical datasets, and only evaluate missingness under MCAR (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008). Further, most proposed methods only perform direct evaluation, using metrics such as MAE, MSE, RMSE, and AUC (Hasan et al., 2021; Joel and Doorsamy, 2022). Very few studies use both direct and indirect evaluation, and even studies that include indirect evaluation only evaluate predictive accuracy (Lin and Tsai, 2020). This leaves a significant research gap, specially for desiderata such as stability and self-consistency, which have recently been shown to be highly sensitive to data-centric interventions (Black and Fredrikson, 2021; Black et al., 2022; Cooper et al., 2023), and unfairness, which can be worsened by improperly chosen data cleaning techniques (Guha et al., 2024). Additionally, the performance of MVI techniques under multi-mechanism missingness (Zhou et al., 2024) and missingness shifts (Zhou et al., 2023) remains unexplored, despite these conditions being more likely to occur in practice due to distribution shifts in production deployments (Graham, 2009).
Notably, and worryingly, overwhelming evidence in the literature indicates that there is no single “best-performing” MVI approach on accuracy  (Hasan et al., 2021; Lin and Tsai, 2020; Shadbahr et al., 2023; García Laencina et al., 2010; Farhangfar et al., 2008), and that model performance (narrowly measured based on ‘correctness’ thus far) depends on dataset characteristics such as size and correlation between variables (Alabadla et al., 2022) and missingness rates in the train and test sets (Lin and Tsai, 2020; Shadbahr et al., 2023).
@@
There has been some recent interest in studying the social harm that can come from poorly chosen MVI techniques (Caton et al., 2022; Zhang and Long, 2022, 2021a, 2021b; Jeong et al., 2022; Wang and Singh, 2021; Martínez-Plumed et al., 2019; Feng et al., 2024). Most empirical studies (Wang and Singh, 2021; Zhang and Long, 2021b, 2022; Caton et al., 2022) have worked with the COMPAS (Larson et al., 2016) and Adult  [^2] https://archive.ics.uci.edu/dataset/2/adult datasets, the latter of which has been “retired” from community use due to issues with provenance (Ding et al., 2021). Further, these experimental studies employ randomly-generated missingness: usually by randomly sampling or using a fixed set of columns, and randomly picking rows in which to replace values with null. We critique this approach, since detecting and mitigating unfairness requires broader socio-technical thinking, such as having higher rates of missingness for minority groups and in features that are highly correlated with sensitive attributes (called proxy variables in the fairness literature) (Caton et al., 2022).
@@
Zhang and Long (2021b) evaluate MVI methods on imputation fairness, defined as the imputation accuracy difference on privileged and disadvantaged demographic groups. They find that (1) higher missingness disparity leads to higher imputation unfairness; (2) higher rates of missingness lead to higher imputation unfairness; (3) more imbalance in the data (or larger disparity in the sample sizes across groups) leads to higher imputation unfairness; (4) the same imputation method has different prediction fairness under different missingness mechanisms. Their evaluation is limited to randomly generated null values in COMPAS. We extend their analysis of imputation fairness to additional datasets, missingness scenarios and imputation fairness definitions (such as KL divergence difference).
Zhang and Long (2022) introduce the notion of imputation fairness risk and provide bounds for “correctly specified” imputation methods. While there is very little theoretical work in this space, and this is a commendable contribution, we call into question the broader implications of this result: it is well-known that imputation quality metrics do not fully capture downstream model performance (Zhang and Long, 2021a). Put differently, a classifier built on poor imputation quality can still give good downstream performance (Shadbahr et al., 2023). With this in mind, we ask the question: Does minimizing imputation unfairness minimize model unfairness? Our empirical findings indicate it does not, as is discussed in Section 4.4.
Missingness and stability. We are not aware of any work investigating the effect of missing value imputation on model stability.
@@ 3.1. Methodology for Simulating Missingness
We start with datasets in which there are no null values, and then simulate missingness. We make this choice because we are interested in comparing MVI performance under single-mechanism versus multi-mechanism missingness, and under missingness shifts, and, to the best of our knowledge, there are no datasets with naturally-occurring documented missingness of this form.
Our methodology for simulating missingness is based on evaluation scenarios, defined by the missingness mechanism during training and testing, shown in Table 1: (1) single-mechanism missingness, injected similarly into train and test sets (S1 - S3); (2) single-mechanism missingness, injected differently into train and test sets (missingness shift) (S4 - S9); and (3) missingness is mixed, to include all three missingness mechanisms, and is injected similarly into train and test sets (S10).
@@
For each dataset in our study (see Section 3.5), we designed socially-salient missingness scenarios corresponding to the three missingness mechanisms (MCAR, MAR, MNAR). Following (Joel and Doorsamy, 2022; Martínez-Plumed et al., 2019), we selected columns in which to inject null values based on the correlation coefficients between all covariates and the target variable, as well as feature importance computed using scikit-learn’s built-in functionality. We used sensitive attributes in the design of MAR missingness to simulate pre-existing bias, as described in Section 1, and injected higher rates of missingness for disadvantaged groups wherever possible (in some cases there were too few samples from disadvantaged groups), aiming to reflect realistic instances of disparate non-response caused by disparate access, distrust, or procedural injustice (Akpinar et al., 2024). Furthermore, we selected columns that may plausibly contain missingness. For example, blood pressure or cholesterol in diabetes is expected to always be observed, but family history or physical activity might be omitted or withheld.
Table 3 presents missingness scenarios for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2. Details for other datasets are available in Appendix B.2.
@@ 3.2. Missing Value Imputation (MVI) Techniques
As discussed in Section 2, there is a large number of competitive MVI techniques in the literature. For our study, we picked 8 techniques, from 5 broad categories in the taxonomy of  Emmanuel et al. (2021), namely: (1) deletion; (2) statistical techniques: median-mode and median-dummy; (3) machine learning-based techniques: miss-forest (Stekhoven and Bühlmann, 2011) and clustering (Gajawada and Toshniwal, 2012); (4) deep learning-based techniques: datawig (Biessmann et al., 2019) and auto-ml (Jäger et al., 2021); and (5) multiple imputation: boostclean (Krishnan et al., 2017). Additional details about these methods are deferred to Appendix A.1.